### v1

# ToolFlow: Boosting LLM Tool-Calling Through Natural and Coherent Dialogue Synthesis

Abstract
Supervised fine-tuning (SFT) is a common method to enhance the tool calling capabilities of Large Language Models (LLMs), with the training data often being synthesized. The current data synthesis process generally involves sampling a set of tools, formulating a requirement based on these tools, and generating the call statements. However, tools sampled randomly lack relevance, making them difficult to combine and thus reducing the diversity of the data. Additionally, current work overlooks the coherence between turns of dialogues, leading to a gap between the synthesized data and real-world scenarios. To address these issues, we propose a Graph-based Sampling strategy to sample more relevant tool combinations, and a Planned-generation strategy to create plans that guide the synthesis of coherent dialogues. We integrate these two strategies and enable multiple agents to synthesize the dialogue data interactively, resulting in our tool-calling data synthesis pipeline ToolFlow. Data quality assessments demonstrate improvements in the naturalness and coherence of our synthesized dialogues. Finally, we apply SFT on LLaMA-3.1-8B using 8,000 synthetic dialogues generated with ToolFlow. Results show that the model achieves tool-calling performance comparable to or even surpassing GPT-4, while maintaining strong general capabilities.
[…]
ToolFlow: Boosting LLM Tool-Calling
Through Natural and Coherent Dialogue Synthesis
Zezhong Wang1thanks: Work done during the internship at Huawei Noah’s Ark Lab., Xingshan Zeng2, Weiwen Liu2, Liangyou Li2, Yasheng Wang2, Lifeng Shang2, Xin Jiang2, Qun Liu2, Kam-Fai Wong1 1The Chinese University of Hong Kong, 2Huawei Noah’s Ark Lab {zzwang,kfwong}@se.cuhk.edu.hk {zeng.xingshan,liuweiwen8,liliangyou}@huawei.com {wangyasheng,Shang.Lifeng,Jiang.Xin,qun.liu}@huawei.com
1 Introduction
Enabling Large Language Models (LLMs) to perform tool calling significantly enhances their capabilities and practical applications. This requires the models to possess strong understanding, reasoning, and instruction-following abilities. Customized fine-tuning is a widely used method to improve the tool-calling capabilities of LLMs Abdelaziz et al. (2024); Patil et al. (2023); Schick et al. (2023); Qin et al. (2023). However, access to fine-tuning data can be limited. One viable solution is to utilize LLMs for data synthesis Basu et al. (2024); Wang et al. (2023); Xu et al. (2023); Yu et al. (2024).
A typical tool-calling data synthesis process involves three steps: (1) selecting candidate tool(s), (2) generating requirements based on those tools, and (3) creating the call statements Tang et al. (2023); Liu et al. (2024b). However, the data synthesized through this method often lacks realism and naturalness. Randomly sampled tools frequently fail to interconnect, making it difficult to combine them for complex tasks. Consequently, the requirements for subsequent synthesis tend to be simplistic, which reduces the diversity and complexity of the data. Furthermore, much of the existing research focuses solely on generating single-turn tool-calling instructions, neglecting the coherence between dialogue turns Qin et al. (2023); Yang et al. (2023). In real-world interactions, LLMs typically engage with users through dialogues rather than single-round Q&A sessions. This creates a gap between Q&A-type training data and its practical application, ultimately diminishing the naturalness of the synthesized data.
To address these two challenges, we propose ToolFlow, a tool-calling data synthesis pipeline that employs a graph-based sampling algorithm to improve the correlation among the selected tools and a planned-generation strategy to enhance the naturalness and coherence of the synthesized tool call dialogues.
Specifically, we consider tools with similar parameters or return values to be related. For instance, both "book_flight" and "get_weather" require parameters related to location. In practical scenarios, these two tools are indeed interconnected, as they often occur together in travel contexts. Based on this assumption, we construct a tool graph that represents the similarity between parameters and return values of the tools. Each node in the graph represents a tool, while the edges indicate the relevance between pairs of tools. When sampling tools, we randomly select a subgraph from this tool graph, ensuring that the sampled tools are more likely to interact effectively, thereby facilitating the generation of complex requirements.
On the other hand, before synthesizing dialogues, we first have the LLM create a plan based on the selected subset of tools. This plan outlines the requests that users need to make in each turn of the dialogue. While constructing the plan, the model focuses on establishing the dialogue framework without worrying about phrasing and details. This approach allows the model to concentrate on the logical relationships and interactions between requirements, resulting in more coherent demands. Additionally, we enable the LLM to incorporate non-tool-call requests into the plan. This not only enhances the diversity of the conversation content but also facilitates seamless transitions between topics, naturally leading to new requirments.
We generate dialogues using three agents: User, Assistant, and Tool. Based on the selected tool subset and the established plan, these agents interact to complete the dialogue. By iterating through the "sampling-planning-generation" process, we synthesized a total of 8,000 dialogues. To evaluate the effectiveness of our proposed method, we conduct a comprehensive ablation study on the graph-based sampling and planning strategy by generating dialogues of the same size selectively without these modules. We perform a thorough evaluation of the data quality, which demonstrates that ToolFlow can effectively enhance the naturalness, coherence, and diversity of the generated dialogues. Finally, we apply supervised fine-tuning to LLaMA-3.1-8B-Instruct Dubey et al. (2024) using the synthesized data and validate improvements in the model’s tool-calling capabilities while preserving its general abilities, with ToolFlow.
We summarize our contributions into the following three key points:
  • We propose a Graph-based Sampling strategy to select related tools, aiming to enhance the diversity and complexity of synthetic tool calling requirements.
  • We introduce a Planned-Generation strategy to improve the naturalness and coherence of synthetic dialogues.
  • We integrate these two strategies and propose ToolFlow, a tool calling dialogue data synthesis pipeline, with extensive experiments and analyses showing its effectiveness.
2 Related Works
Integrating external tools with large language models (LLMs) significantly broadens their functional scope, allowing for more specialized, precise, and reliable solutions to complex problems Qin et al. (2023). There are generally two main strategies for embedding tool-use capabilities into LLMs: prompt-based methods and tool-augmented SFT. Prompt-based methods enable LLMs to utilize tools by providing descriptions and examples of the tools in the prompt, without any incremental training Ruan et al. (2023); Hsieh et al. (2023); Mialon et al. (2023). ReAct Yao et al. (2023) is one of the notable methods within this category. It allows LLMs to switch between reasoning and executing actions to tackle challenging tasks. However, the effectiveness of these approaches can be limited by the inherent capabilities of the model. On the other hand, tool-augmented tuning is attracting increasing interest due to its direct enhancement of LLM’s tool usage capabilities Abdelaziz et al. (2024); Patil et al. (2023); Schick et al. (2023); Qin et al. (2023). As the limited availability of tool calling datasets,  Basu et al. (2024) adapted data from various other domains for application in tool calling studies. Others Liu et al. (2024b); Tang et al. (2023); Qin et al. (2023) mainly synthesized single-turn instructions that involve basic tool calling requirements. However, LLMs typically interact with users through dialogue rather than single-turn Q&A. This mismatch means the data is unnatural, creating a gap with real-world scenarios. Our ToolFlow focuses on enhancing the coherence and naturalness of dialogues in data synthesis, making it more aligned with actual applications.
3 Methodology
Figure 1: The pipeline of dialogue synthesis. The left side shows the Tool Graph with blue boxes representing tools and purple boxes representing parameters or return values. In the middle is the dialogue synthesis plan generated according to sampled tools. On the right is an example of data synthesis by the User, Assistant, and Tool agents.
To generate realistic and coherent dialogues, we propose a three-step data synthesis process: (1) Selecting a tool subset using graph-based sampling; (2) Generating a dialogue plan based on the selected tool subset; (3) Synthesizing dialogues guided by the tool subset and dialogue plan.
3.1 Graph-based Sampling for Tool Selection
Tool calling data synthesis generally starts by selecting one or more tools from the available toolset. While much of the previous work overlooks the significance of tool selection Qin et al. (2023); Patil et al. (2023), relying solely on random sampling, the chosen tools play a crucial role in shaping the quality of the synthesized dialogues. In real-world scenarios, user requirements often necessitate the combined use of multiple tools to achieve a solution. To synthesize more complex user needs, it is vital to ensure that the selected tools can work together. To address this, we propose a Graph Sampling strategy to identify relevant and compatible tool combinations.
3.1.1 Tool Graph Construction
We first construct a graph, $G=(V,E)$, where node $v_{i}\in V$ represents the tool $i$, and the edge $e_{i,j}\in E$ represents whether tool $i$ and tool $j$ are related. The left side of Figure 1 shows an example of the tool graph. We consider tools with similar parameters or return values to be related to each other:
P-P Similarity: When two tools share similar parameters, there is a high probability that the tools are related. For instance, based on "location" and "destination", two semantically similar parameters, we can identify tools like "get_weather" and "book_flight", which are frequently used together in travel-related contexts.
P-R Similarity: If the return value of one tool is similar to the input parameters of another, there is also a high likelihood that the two tools are related. For example, the "check_calendar" tool typically returns the location of events, while the "navigate" tool requires a location as input. When a user requests to "navigate to the location of this afternoon’s meeting," both tools would be called.
To derive similarity between parameters or return values, we first concatenate the name and description of a parameter or a return value using the template "{Name}: {Description}". For example, the parameter “Date” of one specific tool is represented as "Date: Departure date, format as dd/mm/yyyy." Then, we encode these strings using Sentence-BERT Reimers and Gurevych (2019) to obtain the corresponding embeddings. We use $p^{i}_{k}$ and $r^{i}_{l}$ to denote the $k$-th parameters and $l$-th return values of the tool $v_{i}$, respectively. And we use $\mathbf{p}^{i}_{k}$ and $\mathbf{r}^{i}_{k}$ to represent the embeddings of $p^{i}_{k}$ and $r^{i}_{l}$, respectively. The similarity between $v_{i}$’s parameter $p^{i}_{k}$ and $v_{j}$’s parameter $p^{j}_{l}$ is defined as:
$\cos(\mathbf{p}^{i}_{k},\mathbf{p}^{j}_{l})=\frac{\mathbf{p}^{i}_{k}\cdot\mathbf{p}^{j}_{l}}{\|\mathbf{p}^{i}_{k}\|\|\mathbf{p}^{j}_{l}\|}$ (1)
Similarly, the similarity between $v_{i}$’s return value $r^{i}_{k}$ and $v_{j}$’s parameter $p^{j}_{l}$ is defined as $\cos(\mathbf{r}^{i}_{k},\mathbf{p}^{j}_{l})$. If the similarity is greater than a predefined threshold $\tau$, we consider the two tools to be correlated. We set $\tau$ to be 0.82 according to our preliminary study. This means that when the similarity between any pair of parameters from two tools exceeds $\tau$, or when the similarity between a return value and an input parameter of two tools exceeds $\tau$, we assign an edge between the two tools:
$e_{i,j}=\begin{cases}1,&\exists p^{i}_{k}\subseteq v_{i},p^{j}_{l}\subseteq v_{j}:\cos(\mathbf{p}^{i}_{k},\mathbf{p}^{j}_{l})>\tau\\ &\quad\quad\quad\quad\quad\text{or}\\ &\exists r^{i}_{k}\subseteq v_{i},p^{j}_{l}\subseteq v_{j}:\cos(\mathbf{r}^{i}_{k},\mathbf{p}^{j}_{l})>\tau\\ 0,&\text{otherwise}\end{cases}$ (2)
where $i,j=1\cdots N$ and $i\neq j$. We use $\subseteq$ to represent a parameter or a return value is included in the tool.
3.1.2 Graph-based Sampling
With the constructed tool graph, we are able to sample a subset consisting of $n$ tools that might be correlated. Generally, we randomly select a node as the starting point on the graph, and then perform a random walk along the edges of the graph. We stop when the path length reaches $n$, and the nodes included in the path constitute the sampled subset of tools. Details are shown in Algorithm 1.
Algorithm 1 Graph-based Sampling
0:  $G=(V,E)$ with $|V|=N$, integer $n\leq N$ (desired sample size)
0:  Subset $\hat{V}\subseteq V$ with $|\hat{V}|=n$
1:  // Randomly choose a node from $V$ and add it to $\hat{V}$
2:  $\hat{V}\leftarrow\{\text{Uniform}(V)\}$
3:  while $|\hat{V}|<n$ do
4:     $v_{i}\leftarrow\text{last element of }\hat{V}$
5:     // Find all neighbors of $v_{i}$
6:     $N(v_{i})\leftarrow\{v_{j}\in V\mid e_{i,j}=1,\forall e_{i,j}\in E\}$
7:     // Randomly choose a neighbor of $v_{i}$
8:     $v_{j}\leftarrow\text{Uniform}(N(v_{i}))$
9:     if $v_{j}\notin\hat{V}$ then
10:        // Add $v_{j}$ to $\hat{V}$ if not already included
11:        $\hat{V}\leftarrow\hat{V}\cup\{v_{j}\}$
12:     end if
13:  end while
14:  return  $\hat{V}$
3.2 Dialogue Plan Generation
Coherent dialogues usually involve complex tool-calling scenarios, such as cases where the current tool calling relies on the return value of a previous one or where parameters for the current tool calling are already present in the dialogue history. Moreover, realistic dialogues between humans and AI assistants often don’t always require tool use; they frequently involve non-tool-related exchanges, such as chitchat, interspersed with tool-based tasks. To enhance an LLM’s performance in such realistic scenarios, training examples that reflect this balance are essential.
To address the need and enhance the coherence of synthesized dialogues, we propose a Planned-Generation strategy, which indicates planning before generating. We have the LLM first formulates a set of user requirements based on the tool subset to create a dialogue plan. These requirements can involve tool call requests – tasks that necessitate the use of these tools – or non-tool interactions, such as chitchat. The middle part of Figure 1 shows an example of a plan. At this stage, the LLM focuses solely on the logic, coherence, and natural flow of the requirements, without delving into the phrasing of interactions or other nuances. Compared to directly synthesized dialogues, those generated based on a dialogue plan show markedly improved coherence. We provide a detailed assessment of this coherence in the subsequent sections. Please refer to Table 8 for the plan synthetic prompt.
3.3 Multi-Agent Dialogue Synthesis
We set up three agents, user, assistant, and tool, with LLMs to collaboratively synthesize dialogues. The right side of Figure 1 illustrates the synthesis process for one dialogue turn.
The user agent is responsible for initiating requests based on the dialogue plan. It first checks whether the current request in the plan has been completed, which determines whether to continue with the current task or move on to the next one. This ensures that the dialogue stays aligned with the plan’s flow and sequence.
The assistant agent evaluates the user’s request to determine if a tool is required. In cases where no tool is needed, such as chitchat, the assistant agent responds directly. If a tool call is necessary, the assistant verifies whether all required parameters are present based on the tool documentation. If any parameters are missing, the assistant requests clarification from the user; otherwise, it generates the tool call statement.
The tool agent simulates the return values of the requested tool based on the tool documentation and the assistant’s call statement.
The interaction among the three agents continues for each dialogue turn until all the requests in plan are addressed or the preset turn limit is reached. Afterward, all dialogue turns are collected and a rule-based data filtering module is applied to remove low-quality data Liu et al. (2024b, a). The filtering rules primarily check the format of tool call statements, as well as other issues such as incomplete dialogues or missing tool call turns.
3.4 Implementation Details
In this work, for tool selection, we directly utilize the tools from ToolBench Qin et al. (2023) (including over 16,000 RESTful APIs) as our available tools. To standardize tool descriptions, we follow the setting of OpenAI Function Calling [^1] https://platform.openai.com/docs/guides/function-calling and prompt an LLM (Llama-3.1-8B) to convert all these tools into JSON format. For cases where information is incomplete, such as missing parameter descriptions, we rely on the LLM to infer and fill in the missing details during the conversion process. A demonstration tool is presented in Figure 2.
We use GPT-4 OpenAI et al. (2024) for all generative tasks, including dialogue plan generation and agent simulation, unless otherwise specified. Multiple versions of GPT-4 are randomly selected for each dialogue, potentially enhancing diversity.
Setting Statistics
Graph Plan # tokens # call # call turn
8,054,298 21,069 17,112
8,145,545 25,158 21,504
7,956,087 18,117 15,931
8,069,304 23,301 20,804
Table 1: Basic information about the dialogue datasets synthesized by ToolFlow and its ablation settings. # token, # call, and # call turn represent the number of tokens, tool calls, and turns containing tool calls in the dataset, respectively.
4 Data Quality Assessment
4.1 Basic Data Information
In this section, we evaluate the quality of the synthetic data. To assess the effectiveness of the Graph-based Sampling (referred to as Graph) and the Planned-Generation strategy (referred to as Plan), we synthesized three additional sets of comparative data under different conditions: removing Graph, removing Plan, and removing both Graph and Plan. Each dataset contains 8,000 dialogues. Table 1 presents the total number of tokens, the number of tool calls, and the number of dialogue turns containing tool calls for these datasets.
As shown in Table 1, the total number of tokens generated by the different strategies is similar, at approximately 8 million. Dialogues synthesized using the Planned-Generation strategy include more non-tool interactions, resulting in a lower proportion of tool call requests. In contrast, the Graph-based Sampling strategy increases the number of tool calls. This can be attributed to the connections among tools, where relevant information for subsequent tool calls is generally contained in the dialogue history, thereby reducing the need for additional turns to ask for missing information.
4.2 Quality Evaluation
To further assess the quality of the synthetic dialogues, we implemented both an automatic evaluation and a model-based evaluation.
The automatic evaluation primarily assesses the coherence and diversity of the dialogues. Following Dziri et al. (2019), we assess the coherence of the dialogue as a Natural Language Inference (NLI) task. We treat two consecutive turns in the dialogue as the premise and hypothesis, respectively, and calculate the ratio of entailment relation (EnR) as well as the semantic similarity (SS) between them. A higher EnR or SS between turns indicates that the dialogue is more coherent.
Regarding diversity, we calculate the text’s Shannon entropy (H) based on the word frequency Shannon (1948). We also compute the Distinct-N Score Li et al. (2016) for the dataset, with $N=3$ (D-3). Higher entropy or Distinct-N Score indicates that the dataset contains more information and has greater diversity.
In addition, we randomly sampled 200 dialogues in each dataset for the model-based evaluation. We used GPT-4 OpenAI et al. (2024) to carefully evaluate each dialogue based on four dimensions: naturalness (NAT), coherence (COH), helpfulness (HELP), and accuracy (ACC). The prompt for GPT-4 evaluation is shown in Table 9.
The evaluation results are shown in Table 2. There are two key observations:
  • H and D-3 Score demonstrate that Graph Sampling enhances the diversity of the data.
  • Both evaluations (SS, EnR, and COH) show that Planned-generation improves the coherence of the dialogue.
For more detailed settings, analysis and explanations, please refer to Appendix A.1.
Automatic Evaluation
Setting Coherence Diversity
Graph Plan SS EnR H D-3
63.36 47.3 10.36 0.4865
62.03 32.1 10.14 0.4364
61.72 48.1 9.82 0.3393
58.65 35.4 9.75 0.3078
GPT-4 Evaluation
Graph Plan NAT COH HELP ACC
3.72 3.91 4.71 4.92
3.00 3.72 4.51 4.84
3.51 3.88 4.39 4.87
2.93 3.66 4.18 4.90
Table 2: Results of automatic evaluation and GPT-4 evaluation on the data synthesized by ToolFlow and its ablation settings.
5 Experiments
5.1 Settings
Baselines Simple Multiple Parallel Parallel multiple Avg.
Claude-3.5-Sonnet 88.55 95.00 91.50 92.50 91.22
GPT-4-turbo-0409 87.45 96.50 91.00 89.00 90.28
GPT-4o-0513 80.55 91.00 90.00 83.00 85.02
LLaMA-3.1 8B 90.36 89.50 73.50 73.50 83.40
[…]
5.1.1 Datasets
We conducted experiments on the following three tool-calling datasets to validate the tool call capability of the model trained with ToolFlow.
  • BFCL Patil et al. (2023) primarily consists of Python-style tool call data, divided into four categories Simple, Multiple, Parallel, and Parallel Multiple. We selected the categories that can be evaluated with the Abstract Syntax Tree (AST), which are statistically stable and easy to evaluate. The accuracy is reported.
  • API-Bank Li et al. (2023) is a dialogue-style tool call dataset, including two settings: Call and Retrieve + Call. The model is required to call predefined local Python tools based on user requirements in the dialogue. Accuracy is measured by evaluating whether the tool return values match the ground truth.
  • ToolAlpaca Tang et al. (2023) establishes a multi-agent simulator. It utilizes GPT-4 OpenAI et al. (2024) to simulate the return values of tools. The model can make modifications and re-call the tool based on the return values (e.g., error messages). Finally, GPT-4 evaluates the accuracy of Process and Response.
Additionally, to examine changes in general performance, we evaluated the model’s reasoning and conversational abilities using MMLU Hendrycks et al. (2021), BBH Suzgun et al. (2022), and MTBench Zheng et al. (2024).
5.1.2 Models
In our main experiments, we use LLaMA-3.1-8B-Instruct Dubey et al. (2024) as base model to examine the effectiveness of the synthetic dialogues generated with ToolFlow. For simplicity, we use ToolFlow to refer to the fine-tuned model throughout the remainder of this paper. The models we compared include GPT-3.5, GPT-4, GPT-4o OpenAI et al. (2024), Claude Bai et al. (2022), LLaMA-3.1 Dubey et al. (2024), etc, as well as baselines from the paper of the evaluated datasets, such as Lynx-7B Li et al. (2023) and ToolAlpaca-7B Tang et al. (2023), etc. For specific checkpoint information, please refer to the experimental result tables.
5.2 Results
5.2.1 ToolFlow achieves tool-calling capability comparable to GPT-4.
We evaluated ToolFlow’s tool-calling ability on the BFCL. This dataset contains questions from four categories. In the Simple category, each question contains one tool, which the LLM must correctly call based on requirements. The Multiple question includes 2-4 tools, requiring the model to choose and call the most suitable one. In the Parallel category, several tools should be called in one turn. Multiple Parallel adds distracting candidate tools to the Parallel setup.
The results are shown in Tabel 5.1. Compared to the baseline, ToolFlow achieved performance comparable to most mainstream LLMs. ToolFlow slightly outperformed GPT-4 and GPT-4o, but was slightly weaker than Claude-3.5-Sonnet. The difference is mainly reflected in Parallel Multiple type questions. In these types of questions, the model is presented with multiple candidate tools and required to call these tools in parallel, which is of high complexity in nature. We believe that the difference in the model sizes results in ToolFlow’s performance being inferior to Claude, as ToolFlow only has 8B parameters.
The ablation experiment shows that the model trained on data synthesized by strategies including both Graph-based Sampling and Generated Plan performs the best. This is especially evident in Parallel Multiple type questions. Compared to LLaMA-3.1-8B-Instruct, the trained model achieved an accuracy improvement of 11.5% on Parallel Multiple (increasing from 73.5% to 85%).
While different models or training strategies exhibit some variance in certain categories, the differences in average performance are not significant. Therefore, we conducted further comparative experiments on additional tool call datasets.
5.2.2 ToolFlow achieves SOTA on dialogue data.
BFCL tests the model’s tool calling capability in the form of Q&A. However, we believe that a conversational format is closer to real-world application scenarios. On the other hand, our synthesized training data is also in the form of dialogue. Therefore, in the BFCL test, the advantage of ToolFlow cannot be fully demonstrated.
API-Bank is a dialogue dataset. During evaluation, the model needs to make tool call requests after receiving user demands and provides a response based on the tool’s return value. This process may occur multiple times within a single dialogue. It includes two test settings: Call and Retrieve + Call. In the Call setting, the assistant selects tools from a candidate tool set to fulfill user requests. In the Retrieve + Call setting, the assistant only has access to a search tool. The assistant needs to search for the relevant tools first, and then call them.
Baselines L1 L2 Avg.
GPT-4-turbo-0409 72.43 39.26 55.85
GPT-4o-0513 76.19 42.96 59.58
GPT-3.5-turbo-0125 70.43 52.59 61.51
Lynx-7B 49.87 30.37 40.12
Llama-3.1-8B-Instruct 71.18 37.04 54.11
[…]
76.26 42.23 59.25
[…]
76.76 39.27 58.02
Table 4: Results on API-Bank Dataset. L1 and L2 refer to the setting of Call and Retrieve + Call, respectively. Table values are shown as a percentage. indicates results derived from the original paper.
From the results in Table 5.2.2, ToolFlow achieves state-of-the-art average accuracy. Under the Call setting, ToolFlow outperforms all baselines. In the retrieve + call setting, ToolFlow is inferior to GPT-3.5-turbo but superior to other baselines, including GPT-4. The ablation experiments show that Graph-based Sampling strategy can improve the model’s accuracy in tool call under this setting. This is because tools obtained through Graph sampling often have sequential correlations. As a result, the training data includes more examples of sequentially calling tools, aligning better with the requirements of the retrieve + call setting.
Baselines Proc. Resp. Overall
GPT-3.5 77 85 75
ToolAlpaca-13B 63 69 60
ToolAlpaca-7B 70 73 70
LLaMA-3.1 8B 74 80 74
[…]
80 85 80
81 83 79
78 83 77
Table 5: Results on ToolAlpaca Dataset. Proc. and Resp. stand for Procedure and Response, respectively. Table values are shown as a percentage. indicates results derived from the original paper.
5.2.3 ToolFlow can correct mistakes based on error messages.
Correcting errors is a key capability of LLM tool calls Wang et al. (2024). We conducted the tests in the Simulated setting of ToolAlpaca dataset. This dataset established a simulation environment that utilizes GPT-4 to mimic the return values of tools, including the error messages when calls fail. The model is allowed to self-correct based on these error messages and then retry the call. We assess the tool call and correction capabilities of ToolFlow on this dataset.
The dataset evaluates the accuracy of the tool call Procedure and the model’s final Response. The procedure is considered accurate when the model’s call matches the ground truth. The response is considered accurate when the model’s response can satisfy the user’s instruction. If they are both accurate, the model’s Overall performance is considered accurate. This evaluation was conducted by GPT-4. We presented the results in Table 5.2.2.
ToolFlow’s Procedure accuracy reached 85%, surpassing GPT-3.5’s 77%. In Procedure evaluation, correcting errors is considered as redundant actions and therefore judged as incorrect. Hence, this accuracy implies that in most cases, the first tool call of ToolFlow is accurate. On the other hand, ToolFlow’s Response accuracy of 88% is higher than the Procedure accuracy of 85%, indicating that ToolFlow corrected errors in some test cases. This suggests that ToolFlow has the ability to self-correct based on error messages, even though error correction samples are not included in the training data.
5.2.4 ToolFlow’s General Ability Is NOT Compromised by Fine-tuning.
The fine-tuned model risks catastrophic forgetting, where the capability for tool call is enhanced, but other abilities decline. As an AI assistant, LLM’s reasoning and conversational abilities are equally important. Therefore, we tested the tuned model on the MMLU, BBH, and MTBench datasets to examine whether catastrophic forgetting issues have occurred. The results are shown in Table 6.
The test results on MMLU and BBH show that there is no significant difference in performance between the models before and after training. However, on the MTBench dataset, models trained on data without Graph-based Sampling or Plan-Generation exhibited a decline in performance. Notably, in the evaluation of Turn 2, models trained on synthetic data using the Plan-Generation strategy exhibited a slight performance improvement. This improvement is due to Plan-Generation enhancing the naturalness and coherence of synthetic dialogues, thereby boosting the model’s conversational abilities.
Setting MTBench MMLU BBH
Graph Plan Turn 1 Turn 2 Avg.
- - 7.84 6.85 7.34 69.3% 63.2%
8.08 7.16 7.62 69.8% 63.3%
7.39 6.43 6.91 69.3% 63.5%
7.59 7.04 7.31 69.1% 63.2%
7.16 6.85 7.01 68.9% 63.5%
Table 6: Results of general abilities test on MTBench, MMLU, and BBH. The results in the first row are from the model before tuning. The best results are marked in bold. Red indicates a decrease in performance.
6 Correlation Analysis
To further investigate the impact of the diversity and coherence of the dialogue data on model performance, we conducted additional correlation analysis. In the previous experiments, we synthesized a total of $8,000\times 4=32,000$ dialogues by using ToolFlow and its ablation settings. We randomly sampled from these data 10 times, each time selecting 4,000 dialogues to form 10 new training sets. We calculated the diversity metrics D-3 and H, and the coherence metrics SS and EnR for each dataset. Then, we used this data to fine-tune the Llama3.1-8B-Instruct. We tested these ten fine-tuned models on the BFCL and MTBench. Finally, we calculated the Pearson correlation coefficient between the evaluation metrics of the training data and the model performance and reported it in Table 7.
The average results on the BFCL show that both diversity and coherence of the training data contribute a lot to enhancing the model’s tool-calling capabilities. MTBench results show a strong positive correlation between data coherence and the model’s conversational performance, consistent with our assumption. Notably, while we use entropy and Distinct-N scores to assess diversity, the inconsistent correlation between these metrics and model performance suggests they may reflect different dimensions of diversity. On the other hand, coherence does not appear to positively impact the parallel test sets in the BFCL, likely due to the nature of these tests involving multiple calls within a single turn. Nevertheless, while our ToolFlow has demonstrated the benefits of increasing diversity and coherence in Section 4.2, the correlation results in this section further validate their positive effects on overall performance.
Metrics D-3 H SS EnR
BFCL
Simple 0.174 -0.272 0.347 0.118
Parallel 0.250 0.315 -0.168 0.072
Multiple 0.336 -0.042 0.644 0.612
Parallel Multiple -0.274 0.479 -0.310 -0.102
Avg. 0.250 0.273 0.255 0.378
MTBench
Turn 1 -0.087 0.221 0.262 0.298
Turn 2 -0.185 0.045 0.708 0.415
Avg. -0.179 0.146 0.650 0.454
Table 7: The Pearson correlation between evaluation metrics of data and model’s performance. P-values > 0.2 and < 0.2 are marked in green and red, respectively.
7 Conclusion
In this work, we propose Graph-based Sampling and Planned Generation strategies to enhance the diversity and coherence of synthetic data. Based on these two strategies, we introduce a pipeline called ToolFlow for synthesizing tool calling data and generate 8,000 training samples. Using this dataset, we conduct SFT on Llama3.1-8B-Instruct, resulting in improved tool calling capability of the model. Subsequently, we conduct correlation analysis to demonstrate the influence of data diversity and coherence on model performance. This provides a reference for the composition of training data for the tool-enhanced agent.
Limitations
We summarize the limitations in two points.
As described in Section 3.4, the seed data is a pre-collected tool set including 16,000 APIs. Although our ToolFlow can synthesize more diverse data, it is undeniable that the size and diversity of the tool set also affect the diversity of the data. However, how to enrich the seed data has not yet been studied in this work.
On the other hand, ToolFlow utilizes GPT-4 for data synthesis, and then uses this data to train a 8B-model. Therefore, it still falls under the paradigm of using strong models to train weak models. Whether the model can be improved by training on its own synthesized data is still unknown. We believe that this weak-to-strong setting is more challenging but also more meaningful.
Ethic Statement
In this research, GPT-4 was employed as an evaluator and generator in a manner consistent with ethical guidelines. Transparency about its usage, accountability for its outputs, and mitigation of potential biases were prioritized. Data privacy and security were strictly maintained, and the AI’s limitations were acknowledged, ensuring it supplemented rather than replaced human judgment. This approach aimed to enhance the research quality while upholding academic integrity and ethical standards.
Appendix A Appendix
A.1 Details of Data Quality Assessment
Following Dziri et al. (2019), we converted the dialogue data into a Natural Language Inference (NLI) format. In this format, the request and response from the previous dialogue round serve as the premise, and the current round’s request serves as the hypothesis. We then use a trained BERT Reimers and Gurevych (2019) to predict the relationship between the two, and we calculate the ratio of entailment predictions (EnR). A higher proportion indicates greater coherence between consecutive dialogue rounds. Additionally, we measure the semantic similarity (SS) between the premise and hypothesis. We extract sentence representations using BERT and compute their cosine similarity. A higher similarity score indicates a more coherent dialogue.
Regarding diversity, we calculate the text’s Shannon entropy (H) based on the word frequency. We also compute the Distinct-N Score Li et al. (2016) for the dataset, with $N=3$ (D-3). Higher entropy or Distinct-N Score indicates that the dataset contains more information and has greater diversity. In addition, we sampled 200 dialogues in each set. We used GPT-4 to carefully evaluate each dialogue based on four dimensions: naturalness (NAT), coherence (COH), helpfulness (HELP), and accuracy (ACC). The prompt for GPT-4 evaluation is shown in Table 9.
Automatic evaluation indicates that the Planned-Generation strategy enhances conversation coherence. Both coherence metrics, SS and EnR, reflect this improvement. Intuitively, the plan is carefully designed by the model in advance, leading to more coherent dialogue. On the other hand, the Graph Sampling strategy can increase data diversity. This is because the strategy samples tools with strong associations, and the combination of these tools enhances data diversity.
GPT-4’s evaluation indicates that the Planned-Generation strategy enhances the naturalness of dialogue. This metric assesses whether a dialogue could realistically occur in the real world. In data synthesized without a plan, most user requests are tool calls with little chitchat, which is uncommon in real-world scenarios, resulting in lower naturalness. GPT-4’s coherence evaluation closely aligns with automatic assessments. In terms of helpfulness, both Graph Sampling and Planned-Generation strategies show improvement. The low scores in Helpfulness are mainly due to the assistant frequently asking follow-up questions about parameters, which consumes dialogue turns. These strategies help reduce such behavior. The accuracy of synthesized data is high across all four settings, likely due to pre-filtering with quality control tools.
A.2 Prompts and Demonstrations
See Table 8 to 12 for details.
A.3 Train with other base model with ToolFlow
Table 13 and 14 display the results of two other base models fine-tuned with data generated with ToolFlow.
Table 8: Prompt for plan generation
Table 9: Prompt for GPT-4 Data Evaluation
Table 10: Prompt for User agent
Table 11: Prompt for Assistant agent
Table 12: Prompt for Tool agent
Models Simple Multiple Parallel Parallel Multiple Avg.
Mistral-7B-Instruct-v0.1 61.27% 54.00% 50.50% 47.50% 53.32%
+ ToolFlow 86.91% 81.00% 89.00% 80.50% 84.35%
Qwen2-7B 76.73% 63.50% 82.00% 55.50% 69.43%
+ ToolFlow 90.18% 81.50% 89.00% 78.50% 84.80%
Table 13: The results of Mistral 7B and Qwen2 7B on BFCL. ’+ToolFlow’ represents the results of the model after training with our data.
Models L1 L2 Avg.
Mistral-7B-Instruct-v0.1 59.14 32.48 45.81
+ ToolFlow 68.42 45.19 56.81
Qwen2-7B 58.63 25.19 41.91
+ ToolFlow 64.97 36.93 50.95
Table 14: The results of Mistral 7B and Qwen2 7B on the API-Bank dataset. ’+ToolFlow’ represents the results of the model after training with our data.
1{
2 "type": "function",
3 "function": {
4 "name": "getcurrency",
5 "description": "Get the current exchange rate for a specific currency pair",
6 "parameters": {
7 "type": "object",
8 "properties": {
9 "basecurrency": {
10 "type": "string",
11 "description": "The base currency code, e.g., USD"
12 },
13 "targetcurrency": {
14 "type": "string",
15 "description": "The target currency code, e.g., EUR"
16 }
17 },
18 "required": ["basecurrency", "targetcurrency"]
19 },
20 "results": {
21 "type": "object",
22 "properties": {
23 "exchangerate": {
24 "type": "number",
25 "description": "The current exchange rate from base currency to target currency"
26 },
27 "last_updated": {
28 "type": "string",
29 "description": "The date and time when the exchange rate was last updated"
30 }
31 }
32 }
33 }
34}
Figure 2: Example tool in JSON format.